An experiment was conducted to determine whether either firing temperature or furnace position affects the baked density of a carbon anode. The data are as follows:

Firing Temperature

Position

800 (C)

825 (C)

850 (C)

1

570

1063

565

1

565

1080

510

1

583

1043

590

2

528

988

526

2

547

1026

538

2

521

1004

532

Using Fisher’s LSD method, investigate the differences between the mean baked anode density at the three different levels of temperature, which statement is correct?

  • A.

    There is no significant difference between the firing temperatures 800⁰C and 850⁰C

  • B.

    There is no significant difference between the firing temperatures 825⁰C and 800⁰C

  • C.

    There is no significant difference between the firing temperatures 825⁰C and 850⁰C.

  • D.

    There is a significant difference between the firing temperatures 800⁰C and 850⁰C
Answer #1

Applying  Fisher’s LSD method :

critical value of t with 0.05 level and N-k=15 degree of freedom= tN-k= 2.131
Fisher's (LSD) for group i and j =(tN-k)*(sp*√(1/ni+1/nj)   = 36.70
Difference Absolute Value Conclusion
x1-x2 481.67 significant difference
x3-x1 8.83 not significant difference
x3-x2 490.50 significant difference

option A is correct

There is no significant difference between the firing temperatures 800⁰C and 850⁰C